if |z - i| Options: a*) 14 b) 2 c) 28 d) None of the above

if |z - i|
Options: a*) 14
              b)  2
              c)  28
              d)  None of the above
 

Grade:10

5 Answers

Nishant Vora IIT Patna
7 years ago
If |z-i| = ??

PLs complete the question
Divya
7 years ago
I tried posting the question several times, it kept cutting off the rest of the question. Here: 
If | z-1|
 
Options: a*) 14
              b)  2
              c)  28
              d)  None of the above
Divya
7 years ago
It’s been cut off again! 
- then | z + 12 – 6 i | is less than ____.
Options: a*) 14
              b)  2
              c)  28
              d)  None of the above
Divya
7 years ago
Got it! 
[z + 12 – 6 i ] can be rewritten as [ z – i + 12 – 5 i]
=> | z – i |
and => |12 – 5 i | = sqrt ( 12^2 + 5^2) = 13......................(2)
=> | z + 12 – 6 i |
=> | z + 12 – 6 i |
Hence correct answer is option A.
Divya
7 years ago
I don’t know why Askiitians is not able to printout on the screen what is answered in the answer box. Dissappointing.